# Automatic Derivation of an Optimal Task Frame for Learning and Controlling Contact-Rich Tasks

### v1

@@ III-E Average Vector Orientation Frame (AVOF)
Figure 2: Average vector orientation frame for a set of vectors together with a dispersion ellipsoid characterizing the main directions of variation.
@@
1 Function avof($\boldsymbol{c}_{i}$):
       Input: $\boldsymbol{c}_{i}$ with $i=1\dots N$ ;
@@ 1 Function avof($\boldsymbol{c}_{i}$):
III-F Average Screw Intersection Point (ASIP)
The Average Screw Intersection Point (ASIP) is the approximate intersection point for a given set of $N$ unordered screw axes associated with screws $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ with $i=1\dots N$. In [42], a method was proposed to calculate the ASIP for a set of screw twists $\{\boldsymbol{\mathcalboondox{t}}_{i}\}$ by finding the point with minimal average velocity due to all screws. By considering the velocity instead of the distance to the screw axes, the resulting estimate of the ASIP was shown to be more robust to noise. This is because, in the presence of additive noise, the screw axis is not well-defined when the directional component $\boldsymbol{a}_{i}$ is very small. Later, the method was extended to general screws $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ in [31], so that it could be applied to wrenches $\{\boldsymbol{\mathcalboondox{w}}_{i}\}$ in addition to twists $\{\boldsymbol{\mathcalboondox{t}}_{i}\}$.
@@ III-F Average Screw Intersection Point (ASIP)
Figure 3: Average Screw Intersection Point $\prescript{}{}{\boldsymbol{p}}_{asip}$ for a set of screws and its corresponding uncertainty ellipsoid.
@@
1 Function asip($\boldsymbol{\mathcalboondox{s}}_{i}$):
       Input: $\boldsymbol{\mathcalboondox{s}}_{i}$ with $i=1\dots N$ ;
        // set of screws
@@ 1 Function asip($\boldsymbol{\mathcalboondox{s}}_{i}$):
Algorithm 2 Average Screw Intersection Point
@@ IV-A Input Data and Expected Outcome
In general, the task frame $\{tf\}$ can be a freely moving frame. However, this paper limits itself to finding the best possible task frame that has: (1) a fixed origin w.r.t. either a reference frame $\{w\}$ attached to the world $\mathcal{W}$ or w.r.t. a reference frame $\{tl\}$ attached to the tool $\mathcal{TL}$, and (2) a fixed orientation w.r.t. either $\{w\}$ or $\{tl\}$. We refer to this as selecting the viewpoints for the origin and the orientation. Despite these limitations, such an approach is already suitable for a wide variety of practical applications, as abundantly shown in literature and demonstrated in Sections V and VI.
@@ IV-B Conceptual Approach
We hypothesize that the underlying aim of an expert behind the choice of the task frame is to achieve decoupled control according to two principles: (1) decoupling between the directional and moment component of the motion and wrench screws, and (2) decoupling between motion and wrench.
@@ IV-B1 Principle 1: Decoupled Screw Direction and Moment
Decoupled control between the directional and moment components of a screw can be achieved by choosing the reference point for the moment component of the screw to be on the line of the screw axis itself. This reference point is then taken as the origin of the task frame. The advantage is that control corrections in the directional component of the screw performed at this point will not affect the moment component of the screw.
@@ IV-B2 Principle 2: Decoupled Motion and Wrench
IV-C Selection of Origin
Figure 5: Flow chart of the proposed approach. (a) The origin of the task frame is derived. (b) The orientation of the task frame is derived. (c) The demonstration data including pose, twist, and wrench are expressed in the derived task frame.
As explained in Section IV-B origin candidates for the task frame are found by applying the ASIP procedure to both twist and wrench data, and in two ways: to $\boldsymbol{\mathcalboondox{s}}$ and to $\boldsymbol{\widetilde{\mathcalboondox{s}}}$. In addition, to be able to decide about the viewpoint for the origin, we have to perform this procedure for screws that are expressed w.r.t. both the world frame $\{w\}$ and to the tool frame $\{tl\}$. Hence, in total eight applications of the ASIP procedure are required: to ${}_{*}\boldsymbol{\mathcalboondox{s}}$ or ${}_{*}\boldsymbol{\widetilde{\mathcalboondox{s}}}$, where $\boldsymbol{\mathcalboondox{s}}=\boldsymbol{\mathcalboondox{t}}$ or $\boldsymbol{\mathcalboondox{w}}$, and ‘*’ stands for either $w$ or $tl$. This is depicted in boxes a-1 and a-2 of Fig. 5 and yields eight origin candidates $\boldsymbol{p}_{asip}$ with their covariance matrix $\boldsymbol{C}_{asip}$ (box a-3). The eight candidates are reduced to a single origin in three steps. First, for each combination of a screw $\boldsymbol{\mathcalboondox{s}}$ and its mean-subtracted counterpart $\boldsymbol{\widetilde{\mathcalboondox{s}}}$, we select the origin with the least uncertainty, i.e. with the smallest $\operatorname{det}(\boldsymbol{C}_{asip})$. This is depicted in box a-4. Note that the unit of this uncertainty metric is always $[\text{m}^{6}]$ and that this step boils down to selecting between Model 1 or Model 2 for both twist and wrench screws in the two viewpoints. Next, since the uniform metric allows us to combine the results of twist and wrench analyses, the remaining four origin candidates are averaged two by two in each viewpoint, weighted by the inverse of their covariance matrix (box a-5). I.e., given candidates $\boldsymbol{p}_{asip_{t}}$ and $\boldsymbol{p}_{asip_{w}}$ with corresponding $\boldsymbol{C}_{asip_{t}}$ and $\boldsymbol{C}_{asip_{w}}$, the averaged origin and its covariance matrix correspond to:
@@ IV-C Selection of Origin
Finally, to select the least uncertain viewpoint, $\operatorname{det}(\boldsymbol{C}_{avg})$ is again used to decide between the averaged origins in the two viewpoints, as shown in box a-6.
@@ IV-D Selection of Orientation
The AVOF procedure is depicted in boxes b-1 and b-2 of Fig. 5 and yields four orientation candidates $\boldsymbol{U}_{avof}$ with their covariance matrix $\boldsymbol{C}_{avof}$ as found from (17) and (18), respectively (box b-3). These four candidates are reduced to a single orientation in two steps: (1) averaging the two candidates resulting from either $\boldsymbol{\omega}$ or $\boldsymbol{v}$ and either $\boldsymbol{f}$ or $\boldsymbol{m}$ in every viewpoint (box b-4), and (2) selecting the viewpoint with the least resulting uncertainty (box b-5). The first step is further divided into three substeps: (1a) aligning the two orientation candidates, (1b) weighting the covariance matrices, and (1c) calculating the weighted average and its covariance matrix. These substeps are summarized below.
(1a) Aligning two orientation candidates. In every viewpoint, the two AVOF procedures result in two rotation matrices, $\boldsymbol{U}_{1}$ and $\boldsymbol{U}_{2}$, from twist and wrench data, respectively. The purpose is to find the best correspondence between the columns of $\boldsymbol{U}_{1}$ and $\boldsymbol{U}_{2}$ and label the corresponding pairs with $x$, $y$ or $z$ such that we obtain two rotation matrices $\boldsymbol{R}_{1}$ and $\boldsymbol{R}_{2}$ that are more or less aligned, as illustrated in Fig. 6-a, b.
@@ 5       $\boldsymbol{P}(:,c)\leftarrow 0$;
6       $\boldsymbol{P}(r,c)=\operatorname{sign}(\boldsymbol{R}_{\Delta}(j,c))$ ;
@@ 7       $\boldsymbol{R}_{\Delta}(r,:)\leftarrow 0$ ;
Finally, the viewpoint for the orientation is selected in box b-5 by comparing the determinant of the resulting covariance matrix and choosing the viewpoint with the smallest determinant.
@@
IV-E Representation of Demonstration Data in the Task Frame
@@ V Experiments Design
After this validation, the task frame obtained from demonstration data can be utilized to perform a task by a robot. In pursuit of this, a general-purpose controller is proposed based on the derived task frame. The controller’s effectiveness is validated in the various applications. To assess its robustness and generalization capabilities, diverse scenarios involving disturbances or parameter variations are considered. To affirm the significance of utilizing a well-chosen task frame we investigate whether alterations in the task frame result in a deterioration of task performance. Additionally, controlling tasks within the derived task frame serves to validate that the demonstration data (twist, pose and wrench), represented as explained in Section IV-E, is accurate and sufficient, enabling accurate task performance by the robot.
@@ V-A Applications
Figure 7: Setups used for teaching different tasks to the robot: (a) Revolute Joint (b) Peg on Hole Alignment (c) Bottle Opening (d) Prismatic Joint (e) Drawing (f) 2-D Contour Following (g) 3-D Contour Following. The demonstration data for Bottle Opening and 2-D Contour Following are borrowed from [12].
@@ V-A1 Revolute Joint
Interacting with a revolute joint such as when opening a door is a common daily activity. This task has one d.o.f. for motion space and five d.o.f. for wrench space. The tool is equipped with a gripper that can grip one body attached to the joint, in our case a table, and rotate it around the hinge. The setup is shown in Fig. 7-a. The operator demonstrated the task in different positions and with the joint in different horizontal orientations w.r.t. the world frame. The amount of rotation was almost the same in the different trials.
@@ V-A2 Peg on Hole Alignment
This task can be used as an initialization before peg insertion in case it is difficult to provide a good estimate of the hole’s location beforehand. First, we establish a three-point contact between the intentionally misaligned peg and the hole. Then, we rotate the peg until the symmetry axis of the peg is aligned with the symmetry axis of the hole. Unlike a simple revolute joint, this alignment involves a moving rotation axis as shown in [47]. This task has three d.o.f. for both motion and wrench spaces. The peg is mounted on the tool and the hole is attached to the environment as shown in Fig. 7-b. The operator demonstrated the task in the same location and orientation w.r.t. the world frame. However, slight differences in the tool’s orientation among demonstrations are observed due to human variability. The amount of initial misalignment is nearly the same among trials.
@@ V-A3 Bottle Opening
Removing the cap from a bottle is a non-reciprocal task, i.e., we need to apply moment and rotate the bottle opener in the same direction to remove the cap. Hence, unlike the previous tasks, this one requires non-negligible power consumption. An opener is mounted on the tool and a bottle is fixed to the environment as shown in Fig 7-c. At the start of the task, initial contact is well-established, with the opener completely locked. Hence, there are zero and six d.o.f. in the motion and wrench spaces, respectively. During the opening phase, application of the interaction wrench causes motion, mainly but not exclusively in rotation. The demonstration data for this application is borrowed from [12]. The operator demonstrated the task in different positions, with slightly different orientations of the bottle’s symmetry axis w.r.t. the vertical axis of the world frame. The amount of rotation is consistent among trials.
@@ V-A4 Prismatic Joint
Interacting with a prismatic joint such as when opening a drawer is another common daily activity. This task has one d.o.f. for motion space and five d.o.f. for wrench space. The same tool employed for interacting with a Revolute Joint was used, as shown in Fig. 7-d. The operator always demonstrated the task in the same position and orientation w.r.t. the world frame, with almost the same amount of translation.
@@ V-A5 Drawing
Tasks like drawing, writing and pizza cutting with a roller all involve a point-on-plane contact. The objective is to exert a force normal to a surface such as a table and simultaneously translate the tool over the table. In this study we considered drawing. This task has one d.o.f. for wrench space and five d.o.f. for motion space. While the two translational d.o.f. along the table are functional during the demonstrations, the three orientations can be varied freely by the operator. The tool is equipped with a pen-like rod as shown in Fig. 7-e. The operator demonstrated the task in the same location and orientation w.r.t. the world frame. There is not much variation in the tool’s orientation w.r.t. the world frame and the plane of drawing remained the same among the trials.
@@
The objective of this task, shown in Fig. 7-f, is to follow a 2-D contour using a roller at almost constant translational velocity while maintaining an almost constant contact force. Polishing is a typical application. The task has only one d.o.f. for wrench space, while there are five d.o.f. for motion space. One translational d.o.f. along the contour allows us to follow it. Since it is a 2-d.o.f. application, the tool is expected not to move in the two out-of-plane rotational directions and the out-of-plane translational direction, hence maintaining a constant orientation or position in those directions. Finally, the intention was to maintain also a constant in-plane orientation of the tool w.r.t. the world. The tool consists of a rod with a roller at its tip. The demonstration data was again borrowed from [12]. The operator demonstrated the task in different locations but with approximately the same orientation of the workpiece w.r.t. the world frame. Hence, also the plane of the contour remained the same among the trials.
@@
The objective is to follow a 3-D contour with almost constant translational velocity while simultaneously exerting an almost constant amount of force normal to the contour’s edge. The tool is shown in Fig. 7-g, see [31] for more details. It has five wheels which are expected to be in contact during the experiment. The loss of contact for any of these wheels is considered a task failure. The task involves one instantaneous d.o.f. in motion space, i.e. translation along the contour’s edge, and five instantaneous d.o.f. in wrench space, i.e. normal forces on two sides of the contour and three moments. Compared to the previous tasks, this task required much more skill from the operator to complete successful demonstrations, as it was not evident to maintain contact at all wheels during the entire trials. The operator demonstrated the task with the workpiece in the same position and orientation w.r.t. the world frame.

### v3

@@ III-E Average Vector Orientation Frame (AVOF)
Figure 3: Average vector orientation frame for a set of vectors $\{\boldsymbol{c}_{i}\}$ together with a dispersion ellipsoid characterizing the main directions of variation.
@@
1 Function avof( $\{\boldsymbol{c}_{i}\}$):
       Input: $\{\boldsymbol{c}_{i}\}$ with $i=1\dots N$ ;
@@ 1 Function avof( $\{\boldsymbol{c}_{i}\}$):
III-F Average Screw-axes Intersection Point (ASIP)
The Average Screw-axes Intersection Point (ASIP) is the approximate intersection point for a given set of $N$ unordered screw axes associated with screws $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ with $i=1\dots N$. In [42], a method was proposed to calculate the ASIP for a set of screw twists $\{\boldsymbol{\mathcalboondox{t}}_{i}\}$ by finding the point with minimal average velocity due to all screws. By considering the velocity instead of the distance to the screw axes, the resulting estimate of the ASIP was shown to be more robust to noise. This is because, in the presence of additive noise, the screw axis is not well-defined when the directional component $\boldsymbol{a}_{i}$ is very small. Later, the method was extended to general screws $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ in [41], so that it could be applied to wrenches $\{\boldsymbol{\mathcalboondox{w}}_{i}\}$ in addition to twists $\{\boldsymbol{\mathcalboondox{t}}_{i}\}$.
@@ III-F Average Screw-axes Intersection Point (ASIP)
Figure 4: Average Screw-axes intersection point $\prescript{}{}{\boldsymbol{p}}_{asip}$ for a set of screw axes $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ and its corresponding uncertainty ellipsoid.
@@
1 Function asip( $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$):
       Input: $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$ with $i=1\dots N$ ;
        // set of screw axes
@@ 1 Function asip( $\{\boldsymbol{\mathcalboondox{s}}_{i}\}$):
Algorithm 2 Average Screw-axes Intersection Point
@@ IV-A Input Data and Expected Outcome
In general, the task frame $\{tf\}$ can be a freely moving frame. However, this paper limits itself to finding the best possible task frame that has: (1) a fixed origin w.r.t. either a reference frame $\{w\}$ attached to the world $\mathcal{W}$ or w.r.t. a reference frame $\{tl\}$ attached to the tool $\mathcal{TL}$, and (2) a fixed orientation w.r.t. either $\{w\}$ or $\{tl\}$. We refer to this as selecting the viewpoints, $\mathcal{W}$ or $\mathcal{TL}$, for the origin and the orientation. Despite these limitations, such an approach is already suitable for a wide variety of practical applications, as abundantly shown in literature and demonstrated in Sections V and VI.
@@ IV-B Conceptual Approach
Inspired by the related work reviewed in Section II-C, we hypothesize that the underlying aim of an expert behind the choice of the task frame is to achieve decoupled control according to two principles: (1) decoupling between the directional and moment component of the motion and wrench screws, and (2) decoupling between motion and wrench.
@@ IV-B1 Principle 1: Decoupled Screw Direction and Moment
Decoupled control between the directional component $\boldsymbol{a}$ and moment component $\boldsymbol{b}^{o}$ of a screw, where the screw is either a twist or a wrench, can be achieved by choosing the reference point for the moment component of the screw to be on the line of the screw axis itself. This reference point is then taken as the origin of the task frame. The advantage is that control corrections in the directional component of the screw performed at this point will not affect the moment component of the screw.
@@ IV-B2 Principle 2: Decoupled Motion and Wrench
IV-C Selection of Origin (Fig. 6a)
Figure 6: Flow chart of the proposed approach. In this figure, unc. stands for uncertainty. (a) The origin of the task frame is derived, explained in section IV-C. (b) The orientation of the task frame is derived, detailed in section IV-D. (c) The demonstration data including pose, twist, and wrench are expressed in the derived task frame as elaborated in section IV-E.
As explained in Section IV-B origin candidates for the task frame are found by applying the ASIP procedure to both twist and wrench data, and in two ways: to $\boldsymbol{\mathcalboondox{s}}$ and to $\boldsymbol{\widetilde{\mathcalboondox{s}}}$. In addition, to be able to decide about the viewpoint for the origin, we have to perform this procedure for screws that are expressed w.r.t. both the world frame $\{w\}$ and to the tool frame $\{tl\}$. Hence, in total eight applications of the ASIP procedure are required: to ${}_{*}\boldsymbol{\mathcalboondox{s}}$ or ${}_{*}\boldsymbol{\widetilde{\mathcalboondox{s}}}$, where $\boldsymbol{\mathcalboondox{s}}=\boldsymbol{\mathcalboondox{t}}$ or $\boldsymbol{\mathcalboondox{w}}$, and ‘*’ stands for either $w$ or $tl$. This yields eight origin candidates $\boldsymbol{p}_{asip}$ with their covariance matrix $\boldsymbol{C}_{asip}$. The eight candidates are reduced to a single origin in three steps. First, for each combination of a screw $\boldsymbol{\mathcalboondox{s}}$ and its mean-subtracted counterpart $\boldsymbol{\widetilde{\mathcalboondox{s}}}$, we select the origin with the least uncertainty, i.e. with the smallest $\operatorname{det}(\boldsymbol{C}_{asip})$. Note that the unit of this uncertainty metric is always $[\text{m}^{6}]$ and that this step boils down to selecting between Model 1 or Model 2 for both twist and wrench screws in the two viewpoints. Next, since the uniform metric allows us to combine the results of twist and wrench analyses, the remaining four origin candidates are averaged two by two in each viewpoint, weighted by the inverse of their covariance matrix. I.e., given candidates $\boldsymbol{p}_{asip_{t}}$ and $\boldsymbol{p}_{asip_{w}}$ with corresponding $\boldsymbol{C}_{asip_{t}}$ and $\boldsymbol{C}_{asip_{w}}$, the averaged origin and its covariance matrix correspond to:
@@
Finally, to select the least uncertain viewpoint, $\operatorname{det}(\boldsymbol{C}_{avg})$ is again used to decide between the averaged origins in the two viewpoints.
@@
The AVOF procedure yields four orientation candidates $\boldsymbol{U}_{avof}$ with their covariance matrix $\boldsymbol{C}_{avof}$ as found from (18) and (19), respectively. These four candidates are reduced to a single orientation in two steps: (1) averaging the two candidates resulting from either $\boldsymbol{\omega}$ or $\boldsymbol{v}$ and either $\boldsymbol{f}$ or $\boldsymbol{m}$ in every viewpoint and (2) selecting the viewpoint with the least resulting uncertainty. The first step is further divided into three substeps: (1a) aligning the two orientation candidates, (1b) weighting the covariance matrices, and (1c) calculating the weighted average and its covariance matrix. These substeps are summarized below.
(1a) Aligning two orientation candidates. In every viewpoint, the two AVOF procedures result in two rotation matrices, $\boldsymbol{U}_{1}$ and $\boldsymbol{U}_{2}$, from twist and wrench data, respectively. The purpose is to find the best correspondence between the columns of $\boldsymbol{U}_{1}$ and $\boldsymbol{U}_{2}$ and label the corresponding pairs with $x$, $y$ or $z$ such that we obtain two rotation matrices $\boldsymbol{R}_{1}$ and $\boldsymbol{R}_{2}$ that are more or less aligned, as illustrated in Fig. 7a, 7b.
@@ 5       $\boldsymbol{P}(:,c)\leftarrow 0$;
6       $\boldsymbol{P}(r,c)=\operatorname{sign}(\boldsymbol{R}_{\Delta}(r,c))$ ;
@@ 7       $\boldsymbol{R}_{\Delta}(r,:)\leftarrow 0$ ;
Finally, the viewpoint for the orientation is selected by comparing the determinant of the resulting covariance matrix and choosing the viewpoint with the smallest determinant.
@@
IV-E Representation of Data in the Task Frame (Fig. 6c)
@@ V Experiments Design
To validate the benefits of using a well-chosen and parameterized task model, the task model obtained in Section IV-E is utilized to perform demonstrated tasks by a robot. In pursuit of this, a general-purpose controller is designed based on the derived task frame. The controller’s effectiveness is validated in the various applications. To assess its robustness and generalization capabilities, diverse scenarios involving disturbances or parameter variations are considered. To affirm the significance of utilizing a well-chosen task frame we investigate whether alterations in the task frame result in a deterioration of task performance. Additionally, controlling tasks within the derived task frame serves to validate that the demonstration data (twist, pose and wrench), represented as explained in Section IV-E, is accurate and sufficient, enabling accurate task performance by the robot.
@@ V-A Applications
Figure 8: Setups used for teaching different tasks to the robot: (a) Revolute Joint (b) Peg on Hole Alignment (c) Bottle Opening (d) Prismatic Joint (e) Drawing (f) 2-D Contour Following (g) 3-D Contour Following. The pictures are snapshots of the videos in the supplementary material. The demonstration data for Bottle Opening and 2-D Contour Following are borrowed from [18].
@@ V-A1 Revolute Joint
Interacting with a revolute joint such as when opening a door is a common daily activity. This task has one d.o.f. for motion space and five d.o.f. for wrench space. The tool is equipped with a gripper that can grip one body attached to the joint, in our case a table, and rotate it around the hinge. The setup is shown in Fig. 8a. The operator demonstrated the task in different positions and with the joint in different horizontal orientations w.r.t. the world frame. The amount of rotation was almost the same in the different trials.
@@ V-A2 Peg on Hole Alignment
This task can be used as an initialization before peg insertion in case it is difficult to provide a good estimate of the hole’s location beforehand. First, we establish a three-point contact between the intentionally misaligned peg and the hole. Then, we rotate the peg until the symmetry axis of the peg is aligned with the symmetry axis of the hole. Unlike a simple revolute joint, this alignment involves a moving rotation axis as shown in [47]. This task has three d.o.f. for both motion and wrench spaces. The peg is mounted on the tool and the hole is attached to the environment as shown in Fig. 8b. The operator demonstrated the task in the same location and orientation w.r.t. the world frame. However, slight differences in the tool’s orientation among demonstrations are observed due to human variability. The amount of initial misalignment is nearly the same among trials.
@@ V-A3 Bottle Opening
Removing the cap from a bottle is a non-reciprocal task, i.e., we need to apply moment and rotate the bottle opener in the same direction to remove the cap. Hence, unlike the previous tasks, this one requires non-negligible power consumption. An opener is mounted on the tool and a bottle is fixed to the environment as shown in Fig 8c. At the start of the task, initial contact is well-established, with the opener completely locked. Hence, there are zero and six d.o.f. in the motion and wrench spaces, respectively. During the opening phase, application of the interaction wrench causes motion, mainly but not exclusively in rotation. The demonstration data for this application is borrowed from [18]. The operator demonstrated the task in different positions, with slightly different orientations of the bottle’s symmetry axis w.r.t. the vertical axis of the world frame. The amount of rotation is consistent among trials.
@@ V-A4 Prismatic Joint
Interacting with a prismatic joint such as when opening a drawer is another common daily activity. This task has one d.o.f. for motion space and five d.o.f. for wrench space. The same tool employed for interacting with a Revolute Joint was used, as shown in Fig. 8d. The operator always demonstrated the task in the same position and orientation w.r.t. the world frame, with almost the same amount of translation.
@@ V-A5 Drawing
Tasks like drawing, writing and pizza cutting with a roller all involve a point-on-plane contact. The objective is to exert a force normal to a surface such as a table and simultaneously translate the tool over the table. In this study we considered drawing. This task has one d.o.f. for wrench space and five d.o.f. for motion space. While the two translational d.o.f. along the table are functional during the demonstrations, the three orientations can be varied freely by the operator. The tool is equipped with a pen-like rod as shown in Fig. 8e. The operator demonstrated the task in the same location and orientation w.r.t. the world frame. There is not much variation in the tool’s orientation w.r.t. the world frame and the plane of drawing remained the same among the trials.
@@
The objective of this task, shown in Fig. 8f, is to follow a 2-D contour using a roller at almost constant translational velocity while maintaining an almost constant contact force. Polishing is a typical application. The task has only one d.o.f. for wrench space, while there are five d.o.f. for motion space. One translational d.o.f. along the contour allows us to follow it. Since it is a 2-d.o.f. application, the tool is expected not to move in the two out-of-plane rotational directions and the out-of-plane translational direction, hence maintaining a constant orientation or position in those directions. Finally, the intention was to maintain also a constant in-plane orientation of the tool w.r.t. the world. The tool consists of a rod with a roller at its tip. The demonstration data was again borrowed from [18]. The operator demonstrated the task in different locations but with approximately the same orientation of the workpiece w.r.t. the world frame. Hence, also the plane of the contour remained the same among the trials.
@@
The objective is to follow a 3-D contour with almost constant translational velocity while simultaneously exerting an almost constant amount of force normal to the contour’s edge. The tool is shown in Fig. 8g, see [41] for more details. It has five wheels which are expected to be in contact during the experiment. The loss of contact for any of these wheels is considered a task failure. The task involves one instantaneous d.o.f. in motion space, i.e. translation along the contour’s edge, and five instantaneous d.o.f. in wrench space, i.e. normal forces on two sides of the contour and three moments. Compared to the previous tasks, this task required much more skill from the operator to complete successful demonstrations, as it was not evident to maintain contact at all wheels during the entire trials. The operator demonstrated the task with the workpiece in the same position and orientation w.r.t. the world frame.